OpenStudy (anonymous):

f 2tanx/1-tan^2 x = 1 then x can equal:? there can be more than 1 answer x= 7pi/8 + npi x= 5pi/8 + npi x= 3pi/8 + npi x= pi/8 + npi

OpenStudy (anonymous):

*if

OpenStudy (jhannybean):

think about where tan (x) cannot be defined.

OpenStudy (anonymous):

do you mean where it would not be positive on the unit circle?

OpenStudy (anonymous):

would it be: pi/8 and 7pi/8 then?

OpenStudy (anonymous):

wait, not 7pi but 5pi/8?

OpenStudy (anonymous):

is that right?

OpenStudy (jhannybean):

OR we can solve it this way - \[\large \frac{2\tan(x)}{1-\tan^2(x)}= 1\] the function will not exist when the denominator = 0 , so let's solve for the denominator. \[\large 1-\tan^2(x)= (a-b)^2 = (1-\tan(x))(1+\tan(x))\]

OpenStudy (jhannybean):

\[\large (1-\tan(x))(1+\tan(x)) = 0\] once you find your values for tan(x) you will see that these points create vertical asymptotes for your graph.

OpenStudy (anonymous):

Ok. That makes more sense! Thank you!

OpenStudy (jhannybean):

np :)